OpenStudy (abbles):

Rewrite sin^4x so that it involves only the first powers of cosine.

OpenStudy (abbles):

I have a "calculation" section and a "reason" section. So far, I have: \[\frac{ 1 }{ 2 }(1 - \cos(2x)) (\frac{ 1 }{ 2 }(1-\cos2x))\] Using one of those trig identities... now what? do I distribute?

OpenStudy (faiqraees):

What do they imply by first powers of cosine? Like 1+x+x^2.....?

OpenStudy (agent0smith):

Distribute, then apply the same identity to the remaining cos^2 2x

OpenStudy (agent0smith):

@FaiqRaees they mean no exponents other than 1. Like in terms of cos (2x) and cos (4x) etc

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @agent0smith @FaiqRaees they mean no exponents other than 1. Like in terms of cos (2x) and cos (4x) etc \(\color{#0cbb34}{\text{End of Quote}}\) Oh thanks for clarifying it.

OpenStudy (abbles):

\[(\frac{ 1 }{ 2 } - \frac{ \cos2x }{ 2 })(\frac{ 1 }{ 2 }-\frac{ \cos2x }{ 2 })\] What remaining cos^2(2x) do you mean? Did you mean the cos2x?

OpenStudy (faiqraees):

Now multiply the terms together

OpenStudy (faiqraees):

Upon multiplying the cosine function, you will get a cosine function raised to the power 2. He was talking about that function

OpenStudy (agent0smith):

Yes. But Abbles, make it easier, not harder \[\large \frac{ 1 }{ 2 }(1 - \cos(2x)) (\frac{ 1 }{ 2 }(1-\cos2x)) = \frac{ 1 }{ 4 }(1 - \cos(2x))(1-\cos(2x))\]

OpenStudy (abbles):

\[\frac{ 1 }{ 4 }(1-2\cos(2x) + \cos^2(2x))\] Look about right so far?

OpenStudy (agent0smith):

Yep, now expand the cos^2 term using the same process as earlier.

OpenStudy (abbles):

\[\frac{ 1 }{ 4 }(1 - 2cosx(2x) + \frac{ 1 + \cos4 }{ 2 })\] Oh boy...

OpenStudy (abbles):

Next multiply the 1/4 through?

OpenStudy (abbles):

\[\frac{ 1 - 2cosx(2x) + \frac{ 1+\cos4 }{ 2 }}{ 4 }\]

OpenStudy (agent0smith):

cos 4x but yes. I'd prob just keep the 1/4 outside for neatness.

OpenStudy (agent0smith):

And you have an x(2x)

OpenStudy (abbles):

cos 4x ? hah! didn't see the extra x :P

OpenStudy (abbles):

\[\frac{ 1 - 2\cos(2x) }{ 4 } + \frac{ 1 + \cos4 }{ }\] Could I rewrite it like this?

OpenStudy (agent0smith):

cos 4x not cos 4 ;P \[\large \frac{ 3 }{ 8 }-\frac{1 }{ 2 }\cos 2x+\frac{ 1 }{ 8 }\cos 4x\]

OpenStudy (abbles):

\[\frac{ 3 - 4\cos(2x) + \cos4x }{ 8 }\] Is this right?

OpenStudy (agent0smith):

Yes, same as what i have above.

OpenStudy (abbles):

Okay, thanks - I think this is what they're looking for. For some reason the less doesn't factor out 1/8... is that preferred?

OpenStudy (agent0smith):

It doesn't matter, depends on the book I guess. Yours or mine look fine.

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=sin%5E4+x wolfram gives it with the 1/8 factored but i doubt it matters.

OpenStudy (abbles):

Thanks :)

OpenStudy (agent0smith):

Welcome :)